OpenStudy (anonymous):

2x-6 = (x-3)e^sinx how do you get rid e^sinx? also, what is this type of problem called?

OpenStudy (anonymous):

2x-6/x-3=e^sinx 2(x-3)/x-3=e^sinx 2=e^sinx Ln2=sinx x=arcsin (Ln2)

OpenStudy (anonymous):

it's an equation

OpenStudy (anonymous):

what type of equation? with the base e^sinx

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

not any speial type. You wellcome

OpenStudy (lgbasallote):

2x-6 = (x-3)e^sinx 2x - 6 = xe^sinx - 3e^sinx 3e^sinx - 6 = xe^sinx - 2x 3e^sinx - 6 = x(e^sinx -2) x = \(\frac{3e^{\sin x} - 6}{e^{\sin x} -2}\) that's how you get rid of it @camialtrs93 take note of the method as it is used a lot

OpenStudy (anonymous):

and how you solve for x now?

OpenStudy (anonymous):

@lgbasallote

OpenStudy (lgbasallote):

\[\large x = \frac{3e^{\sin x} - 6}{e^{\sin x} -2}\] \[x = \frac{3\cancel{(e^{\sin x} - 2)}}{\cancel{e^{\sin x} - 2}}\] @myko

OpenStudy (anonymous):

:)

OpenStudy (lgbasallote):

the asker wanted the e^sinx gone lol

OpenStudy (anonymous):

oh alright, thanks @lgbasallote!